OpenStudy (anonymous):

the graph of y=3/x is translated four units up. What is the eqation of the new graph? I am having a hard time with this any help?

OpenStudy (anonymous):

add 4 y=(3/x) + 4

OpenStudy (anonymous):

y = f(x) + a graph of f(x) moves up a units y = f(x) - a graph of f(x) moves down a units

OpenStudy (anonymous):

ok with that - do you need more transformations?

OpenStudy (anonymous):

No thanks!....easier than i was making it!

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks you!!!